# Generalized Entropy Calibration for Inference with Partially Observed Data: A Unified Framework

### v1

@@ 6 Illustrative Examples
We now describe how the proposed framework applies to three important settings. In each case, we specify the decomposition $(O_{i},M_{i})$, the estimating function $U(\theta;Z_{i})$, and the calibration function $b(\theta;O_{i})$; the calibration weights and estimator are then obtained by applying the general procedure in Sections 35.
@@ 6.1 Causal inference
Let $T_{i}\in\{0,1\}$ be a binary treatment indicator, $Y_{i}(1)$ and $Y_{i}(0)$ are the potential outcomes, and $X_{i}\in\mathbb{R}^{p}$ the observed covariates. Under SUTVA, unconfoundedness, and positivity (Rubin, 1974; Rosenbaum and Rubin, 1983; Imbens and Rubin, 2015), the observed outcome is $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$ and the average treatment effect (ATE) is
@@
where $\theta_{t}$ solves $\mathbb{E}\{U_{t}(\theta_{t};X,Y(t))\}=0$ with $U_{t}(\theta_{t};X,Y(t))=Y(t)-\theta_{t}$ for $t\in\{0,1\}$.
@@
$\min_{\omega_{11},\ldots,\omega_{1N}}\sum_{i=1}^{N}T_{i}\,G(\omega_{1i}),$ (6.1)
$\text{subject to}\quad\sum_{i=1}^{N}T_{i}\,\omega_{1i}=N,$ (6.2)
@@ 6.2 Semi-supervised learning
In many applied settings, the outcome variable is expensive or difficult to obtain while covariates are readily available. For example, in electronic health records research, clinical diagnoses require expert review but demographic and laboratory covariates are routinely recorded (Gronsbell and Cai, 2018; Chakrabortty and Cai, 2018). Semi-supervised methods aim to exploit the large pool of unlabeled covariates to improve efficiency over supervised analysis that uses only the labeled sample.
@@
The selection mechanism may be either MAR, with $P(\delta_{i}=1\mid X_{i})=\pi(X_{i};\phi)$ estimated by logistic regression, or MCAR, with $P(\delta_{i}=1)=n/(n+N)$. In both cases, the calibration weights are obtained from (5.2)–(5.4) and the estimator from Algorithm 1. Concretely, the proposed estimator of $\theta$ solves the weighted estimating equation
$\frac{1}{n+N}\sum_{i=1}^{n+N}\delta_{i}\,\omega_{i}\,U(\theta;x_{i},y_{i})=0,$ (6.5)
@@ 6.3 Missing covariates in regression
Suppose $(Y_{i},X_{i})$ with $X_{i}=(X_{1i},X_{2i})$ are related by a regression model $Y_{i}=\mu(X_{1i},X_{2i};\beta)+\epsilon_{i}$ with $\mathbb{E}(\epsilon_{i}\mid X_{i})=0$, and $X_{2i}$ is subject to missingness. Here $O_{i}=(X_{1i},Y_{i})$, $M_{i}=X_{2i}$, and $\delta_{i}=\mathds{1}(X_{2i}\text{ observed})$. Under MAR,
$\delta_{i}\perp X_{2i}\mid(X_{1i},Y_{i}),$
so that the probability of observing $X_{2i}$ may depend on both the fully observed covariates $X_{1i}$ and the response $Y_{i}$, but not on the missing covariate $X_{2i}$ itself. This conditioning on $Y_{i}$ distinguishes the missing-covariate setting from the missing-outcome case: the propensity score model is
@@
$\hat{\phi}=\arg\max_{\phi}\sum_{i=1}^{N}\bigl[\delta_{i}\log\{\pi(X_{1i},Y_{i};\phi)\}+(1-\delta_{i})\log\{1-\pi(X_{1i},Y_{i};\phi)\}\bigr].$
@@
and the calibration function is $b(\beta;X_{1i},Y_{i})=U(\beta;X_{1i},\hat{X}_{2i},Y_{i})$, where $\hat{X}_{2i}$ are cross-fitted predictions of $X_{2i}$ based on $(X_{1i},Y_{i})$. The calibration weights and estimator of $\beta$ are obtained by applying (5.2)–(5.4) and Algorithm 1.
@@ 7 Simulation Study
We evaluate the proposed Hellinger distance (HD) and exponential tilting (ET) entropy calibration estimators across the three application settings. All simulations use $K=4$ folds for cross-fitting unless otherwise noted.
@@ 7.1 Causal inference
Design.
We consider a $2\times 2$ factorial design with two outcome regression models (OR1, OR2) and two propensity score models (PS1, PS2), using $M=500$ Monte Carlo replicates with sample size N=1, [^0] N=1{,}000. Covariates are generated as $\bm{x}_{i}=(x_{i1},x_{i2},x_{i3},x_{i4})^{\top}\sim N(0_{4},I_{4})$, and potential outcomes as $y_{i}(t)=m_{t}(\bm{x}_{i})+\varepsilon_{it}$, $\varepsilon_{it}\sim N(0,1)$, $t\in\{0,1\}$.
  • OR1 (linear): $m_{1}(x)=1+\mathbf{1}^{\top}\bm{x}$, $m_{0}(x)=\mathbf{1}^{\top}\bm{x},\mathbf{1}=(1,1,1,1)^{\top}$; true ATE =1=1.
  • OR2 (nonlinear): Let $h(x)=(x-1)^{3}-x^{2}+x/\{1+\exp(\mathrm{clip}(x))\}+10$ with $\mathrm{clip}(x)=\max\{\min(x,3),-3\}$ applied componentwise. Set $m_{1}(x)=10+\mathbf{1}^{\top}\bm{x}+\tfrac{1}{2}\mathbf{1}^{\top}h(\bm{x})$, $m_{0}(\bm{x})=\mathbf{1}^{\top}\bm{x}+\tfrac{1}{2}\mathbf{1}^{\top}h(\bm{x})$; true ATE =10=10.
  • PS1 (logistic linear): $\pi(\bm{x})=\mathrm{expit}(-0.25+x_{1}+0.5x_{2}-0.5x_{3}-0.1x_{4})$.
  • PS2 (logistic nonlinear): $\pi(\bm{x})=\mathrm{expit}(x_{1}-0.5x_{1}x_{2}-x_{3}^{2}+0.5x_{4}^{3})$.
The treatment indicator is $T_{i}\sim\mathrm{Bernoulli}\{\pi(x_{i})\}$ and the observed outcome is $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$. We compare HD and ET with seven existing methods: inverse probability weighting (IPW; Horvitz and Thompson, 1952), entropy balancing (EBPS; Hainmueller, 2012), optimal covariate balancing (oCBPS; Fan et al., 2022), covariate balancing propensity score (CBPS; Imai and Ratkovic, 2014), empirical balancing calibration weighting (EBCW; Chan et al., 2016), and AIPW with either a linear model or GAM for the outcome regression, denoted AIPW (LM) and AIPW (GAM).
@@
Results.
Figure 1 summarizes the Monte Carlo distributions across four specification regimes. Under OR1PS1 (both models correct), all methods show negligible bias and similar variability, consistent with the local efficiency result in Remark 1: when both models are correctly specified, the proposed estimator matches AIPW.
@@ 7.2 Semi-supervised learning
Design.
The target is the regression coefficient $\beta$ in $y_{i}=x_{i}^{\top}\beta+\epsilon_{i}$. Covariates are $x_{i}\sim N(0_{p},I_{p})$ with $p=4$, total sample size n+N=2,000n+N=2{,}000, and M=1, [^0] M=1{,}000 replications. The outcome is generated under two specifications:
@@
Results.
Under OR model misspecification (OR2) with MAR (Figure 2), HD and ET achieve the highest efficiency among all methods for all five regression coefficients $\beta_{0},\ldots,\beta_{4}$. The improvement is most visible for $\beta_{0}$: the interquartile range of HD and ET is roughly half that of the supervised estimator, and noticeably smaller than those of PI, EASE, DRESS, and PSSE. The existing semi-supervised methods (PI, EASE, DRESS, PSSE) provide only modest gains over the supervised estimator under OR misspecification, because their efficiency improvements rely on the outcome model being approximately correct. In contrast, our proposed estimators benefit from the debiasing constraint, which compensates for outcome-model misspecification through the propensity-score information.
@@ Results.
Figure 2: Boxplots of the estimated linear regression coefficients $\bm{\beta}$ under OR2 (OR model misspecification) with MAR missingness.
Figure 3: Boxplots of the estimated linear regression coefficients $\bm{\beta}$ under OR2 (OR model misspecification) with MCAR missingness.
Figure 4: Estimation of the linear regression coefficients $\bm{\beta}$ under OR1 (when the OR model is correctly specified) and MCAR missingness mechanism.
@@ 7.3 Missing covariates in regression
Design.
We estimate $\beta=(\beta_{0},\beta_{1},\beta_{2})$ in the linear model $y_{i}=\beta_{0}+\beta_{1}x_{1i}+\beta_{2}x_{2i}+\epsilon_{i}$, where $x_{1i}\sim N(0,1)$, $x_{2i}\sim\mathrm{Bernoulli}(0.5)$, $\epsilon_{i}\sim N(0,1)$, and $x_{2i}$ is subject to missingness. Sample size is $N=500$ with M=1, [^0] M=1{,}000 replications. Two PS models govern the missingness of $x_{2i}$: PS1 (MAR), $\mathrm{logit}(\pi_{i})=-1+0.5x_{1i}+0.5y_{i}$; PS2 (MCAR), $\mathrm{logit}(\pi_{i})=-1$. Two OR models generate $y_{i}$:
@@
Results.
Table 2 reports bias, standard error, and RMSE ($\times 10$) across all scenarios. The full-sample estimator serves as a benchmark with negligible bias and smallest RMSE. The CC estimator shows substantial bias under MAR (OR1PS1 and OR2PS1), confirming that discarding incomplete cases introduces systematic error when the missingness depends on the observed data.

### v2

@@ 6 Illustrative Examples
We now illustrate how the proposed framework specializes to three important settings: causal inference, semi-supervised learning, and regression with missing covariates. The goal of this section is not merely to present three applications of the same algorithm, but to show that these problems share a common inferential structure. In each case, the target parameter is defined by an estimating equation involving a full-data vector $Z_{i}=(O_{i}^{\top},M_{i}^{\top})^{\top}$, while only $O_{i}$ is always observed. The proposed method proceeds in the same way across all three settings: we identify the observed part $O_{i}$, the missing part $M_{i}$, the estimating function $U(\theta;Z_{i})$, and a calibration function $b(\theta;O_{i})$ approximating $\mathbb{E}\{U(\theta;Z_{i})\mid O_{i}\}$. The calibration weights and the resulting estimator are then obtained by applying the general procedure in Sections 35.
These examples clarify the unifying role of calibration. In causal inference, the missing component is a counterfactual outcome; in semi-supervised learning, it is an unlabeled response; in regression with missing covariates, it is an incompletely observed predictor. Although the substantive interpretations differ, the same calibration logic applies in all three cases. Table 2 summarizes the missing data framework for the partially observed data settings.
Table 2: Common partially observed-data template across the three applications.
Setting $O_{i}$ $M_{i}$ $b(\theta;O_{i})$
Causal inference $X_{i}$ $Y_{i}(t)$ $\hat{Y}_{i}(t)-\theta_{t}$
Semi-supervised learning $X_{i}$ $Y_{i}$ $U(\theta;X_{i},\hat{Y}_{i})$
Missing covariates $(X_{1i},Y_{i})$ $X_{2i}$ $U(\beta;X_{1i},\hat{X}_{2i},Y_{i})$
@@ 6.1 Causal inference
We begin with causal inference, where the missing-data interpretation is especially transparent. For each unit, one potential outcome is observed and the other is missing, so estimation of the average treatment effect can be viewed as inference with a partially observed full-data vector. Let $T_{i}\in\{0,1\}$ denote treatment assignment, let $Y_{i}(1)$ and $Y_{i}(0)$ be the potential outcomes, and let $X_{i}\in\mathbb{R}^{p}$ be pretreatment covariates. Under SUTVA, unconfoundedness, and positivity (Rubin, 1974; Rosenbaum and Rubin, 1983; Imbens and Rubin, 2015), the observed outcome is
$Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0),$
and the average treatment effect is
@@
where $\theta_{t}$ solves $\mathbb{E}\{U_{t}(\theta_{t};X,Y(t))\}=0$ with
$U_{t}(\theta_{t};X,Y(t))=Y(t)-\theta_{t}\,\,\text{for}\,\,t\in\{0,1\}.$
@@
$\min_{\omega_{11},\ldots,\omega_{1N}}\sum_{i=1}^{N}T_{i}\,G(\omega_{1i}),$ (16)
$\text{subject to}\quad\sum_{i=1}^{N}T_{i}\,\omega_{1i}=N,$ (17)
@@ 6.2 Semi-supervised learning
Semi-supervised learning provides a second example in which the missing-data interpretation is immediate. Here covariates are observed for a large unlabeled sample, but outcomes are available only for a labeled subset. From the perspective of the present framework, the unlabeled units are simply observations with missing responses. The target parameter is again defined through a full-data estimating equation, and the main statistical task is to recover that equation efficiently using both the labeled and unlabeled samples.
Let $X_{i}$ denote the covariates and $Y_{i}$ the outcome of interest. Define $\delta_{i}=1$ if $Y_{i}$ is observed and $\delta_{i}=0$ otherwise, so that $O_{i}=X_{i}$ and $M_{i}=Y_{i}$. Under MAR,
$\delta_{i}\perp Y_{i}\mid X_{i},$
the response mechanism can be modeled through $\pi(X_{i};\phi)=P(\delta_{i}=1\mid X_{i})$.
@@
The selection mechanism may be either MAR, with $P(\delta_{i}=1\mid X_{i})=\pi(X_{i};\phi)$ estimated by logistic regression, or MCAR, with $P(\delta_{i}=1)=n/(n+N)$. In both cases, the calibration weights are obtained from (13)–(15) and the estimator from Algorithm 1. Concretely, the proposed estimator of $\theta$ solves the weighted estimating equation
$\frac{1}{n+N}\sum_{i=1}^{n+N}\delta_{i}\,\omega_{i}\,U(\theta;x_{i},y_{i})=0,$ (20)
@@ 6.3 Missing covariates in regression
Our third example concerns regression with missing covariates. Unlike the previous two settings, the outcome is fully observed, while part of the predictor vector is missing. This changes the form of the propensity-score model: because the response indicator refers to covariate observation rather than outcome observation, the probability of observing the missing covariate may depend on the observed response as well as on the fully observed covariates.
Suppose $(Y_{i},X_{i})$ satisfies the regression model
$Y_{i}=\mu(X_{1i},X_{2i};\beta)+\varepsilon_{i},\qquad\mathbb{E}(\varepsilon_{i}\mid X_{i})=0,$
where $X_{i}=(X_{1i},X_{2i})$ and $X_{2i}$ is subject to missingness. Here
$O_{i}=(X_{1i},Y_{i}),\qquad M_{i}=X_{2i},$
and under MAR,
$\delta_{i}\perp X_{2i}\mid(X_{1i},Y_{i}).$
Thus, the probability of observing $X_{2i}$ may depend on both the fully observed covariates $X_{1i}$ and the response $Y_{i}$, but not on the missing covariate $X_{2i}$ itself. This conditioning on $Y_{i}$ distinguishes the missing-covariate setting from the missing-outcome case: the propensity score model is
@@
$\hat{\phi}=\arg\max_{\phi}\sum_{i=1}^{N}\left[\delta_{i}\log\pi_{i}+(1-\delta_{i})\log(1-\pi_{i})\right],$
where $\pi_{i}=\pi(X_{1i},Y_{i};\,\phi).$
@@
and the calibration function is $b(\beta;X_{1i},Y_{i})=U(\beta;X_{1i},\hat{X}_{2i},Y_{i})$, where $\hat{X}_{2i}$ are cross-fitted predictions of $X_{2i}$ based on $(X_{1i},Y_{i})$. The calibration weights and estimator of $\beta$ are obtained by applying (13)–(15) and Algorithm 1.
@@ 7 Simulation Study
This section evaluates the finite-sample behavior of the proposed generalized entropy calibration estimators in the three application settings introduced in Section 6. The goal is not simply to compare a collection of competing methods, but to examine whether the empirical results reflect the main theoretical properties established in Section 4. In particular, the simulations are designed to assess four questions: whether the proposed estimators recover the efficient benchmark when both working models are correctly specified; whether they remain accurate when only one of the propensity-score or outcome-regression models is correct; whether they improve upon classical AIPW under correct PS specification and misspecified OR models, as predicted by Corollary 3; and whether these qualitative patterns persist across causal inference, semi-supervised learning, and regression with missing covariates.
We report results for the Hellinger distance (HD) and exponential tilting (ET) versions of generalized entropy calibration. Unless otherwise noted, all simulations use $K=4$ folds for cross-fitting.
@@ 7.1 Causal inference
We begin with causal inference because it provides the clearest setting for isolating the doubly robust behavior of the proposed estimator. The simulation uses a $2\times 2$ factorial design combining two outcome-regression specifications and two propensity-score specifications. This design creates the four regimes most relevant to the theory: joint correctness, OR-only correctness, PS-only correctness, and joint misspecification. It therefore allows us to examine not only double robustness, but also the comparative efficiency gains predicted by Corollary 3 when the PS model is correct and the OR model is misspecified.
Design. We consider two outcome-regression models (OR1, OR2) and two propensity-score models (PS1, PS2), using $M=500$ Monte Carlo replicates with sample size N=1, [^0] N=1{,}000. Covariates are generated as $x_{i}=(x_{i1},x_{i2},x_{i3},x_{i4})^{\top}\sim N(0_{4},I_{4})$, and potential outcomes are generated as
$y_{i}(t)=m_{t}(x_{i})+\epsilon_{it},\qquad\epsilon_{it}\sim N(0,1),\quad t\in\{0,1\}.$
  • OR1 (linear): $m_{1}(x)=1+\mathbf{1}^{\top}\boldsymbol{x}$, $m_{0}(x)=\mathbf{1}^{\top}\boldsymbol{x},\mathbf{1}=(1,1,1,1)^{\top}$; true ATE =1=1.
  • OR2 (nonlinear): Let $h(x)=(x-1)^{3}-x^{2}+x/\{1+\exp(\mathrm{clip}(x))\}+10$ with
    $\mathrm{clip}(x)=\max\{\min(x,3),-3\}$
    applied componentwise. Set $m_{1}(x)=10+\mathbf{1}^{\top}\boldsymbol{x}+\tfrac{1}{2}\mathbf{1}^{\top}h(\boldsymbol{x})$, $m_{0}(\boldsymbol{x})=\mathbf{1}^{\top}\boldsymbol{x}+\tfrac{1}{2}\mathbf{1}^{\top}h(\boldsymbol{x})$; true ATE =10=10.
  • PS1 (logistic linear): $\pi(\boldsymbol{x})=\mathrm{expit}(-0.25+x_{1}+0.5x_{2}-0.5x_{3}-0.1x_{4})$.
  • PS2 (logistic nonlinear): $\pi(\boldsymbol{x})=\mathrm{expit}(x_{1}-0.5x_{1}x_{2}-x_{3}^{2}+0.5x_{4}^{3})$.
The resulting four combinations of OR and PS models allow us to check whether the proposed calibration estimators match AIPW under joint correctness, remain reliable when only one working model is correct, and exhibit the clearest gains precisely in the PS-correct/OR-misspecified regime emphasized by the theory.
We compare HD and ET with seven existing methods: inverse probability weighting (IPW; Horvitz and Thompson, 1952), entropy balancing (EBPS; Hainmueller, 2012), optimal covariate balancing (oCBPS; Fan et al., 2022), covariate balancing propensity score (CBPS; Imai and Ratkovic, 2014), empirical balancing calibration weighting (EBCW; Chan et al., 2016), and AIPW with either a linear model or GAM for the outcome regression, denoted AIPW (LM) and AIPW (GAM).
@@
Results. Figure 1 summarizes the Monte Carlo distributions across four specification regimes. Under OR1PS1 (both models correct), all methods show negligible bias and similar variability, consistent with the local efficiency result in Remark 1: when both models are correctly specified, the proposed estimator matches AIPW.
@@ 7.2 Semi-supervised learning
We next consider semi-supervised learning to examine whether the same qualitative pattern carries over to a setting in which the missing component is an unlabeled response rather than a counterfactual outcome. From the perspective of the present framework, this is again a partially observed-data problem: covariates are observed for all units, while responses are available only for a labeled subset. The simulation is designed to assess how effectively the proposed calibration estimators use the unlabeled covariates for efficiency, while the debiasing constraint protects against bias when the working prediction model is imperfect or when the labeled and unlabeled samples differ in distribution.
The main question in this experiment is whether the same advantages seen in the causal-inference setting persist when the balancing constraint is constructed from cross-fitted predictions of the missing response. In particular, we examine whether ET and HD continue to recover the efficient benchmark under favorable specification, remain stable under partial misspecification, and improve upon purely supervised or outcome-model-based semi-supervised competitors when the prediction model is imperfect.
Design. The target is the regression coefficient $\beta$ in $y_{i}=x_{i}^{\top}\beta+\epsilon_{i}$. Covariates are $x_{i}\sim N(0_{p},I_{p})$ with $p=4$, total sample size n+N=2,000n+N=2{,}000, and M=1, [^0] M=1{,}000 replications. The outcome is generated under two specifications:
@@
Results. Under OR model misspecification (OR2) with MAR (Figure 2), HD and ET achieve the highest efficiency among all methods for all five regression coefficients $\beta_{0},\ldots,\beta_{4}$. The improvement is most visible for $\beta_{0}$: the interquartile range of HD and ET is roughly half that of the supervised estimator, and noticeably smaller than those of PI, EASE, DRESS, and PSSE. The existing semi-supervised methods (PI, EASE, DRESS, PSSE) provide only modest gains over the supervised estimator under OR misspecification, because their efficiency improvements rely on the outcome model being approximately correct. In contrast, our proposed estimators benefit from the debiasing constraint, which compensates for outcome-model misspecification through the propensity-score information.
@@
Figure 2: Boxplots of the estimated linear regression coefficients $\beta$ under OR2 (OR model misspecification) with MAR missingness.
Figure 3: Boxplots of the estimated linear regression coefficients $\beta$ under OR2 (OR model misspecification) with MCAR missingness.
Figure 4: Estimation of the linear regression coefficients $\beta$ under OR1 (when the OR model is correctly specified) and MCAR missingness mechanism.
@@ 7.3 Missing covariates in regression
Our third simulation considers regression with missing covariates. This setting is conceptually important because the incompletely observed component is now a predictor rather than an outcome, and the response mechanism may depend on the observed response as well as on the fully observed covariates. The purpose of this experiment is to determine whether the same calibration logic continues to deliver the expected robustness and efficiency gains in a structurally different missing-data problem.
As in the previous settings, the comparison is organized around the behavior of the method under correct and misspecified working models. Of particular interest is the regime in which the propensity-score model is correctly specified but the regression-based imputation model for the missing covariate is misspecified. According to Corollary 3, this is the setting in which the debiasing constraint should provide the clearest improvement over classical AIPW by enlarging the effective calibration space. The simulation therefore serves as a direct finite-sample check of whether the variance-dominance phenomenon identified in Section 4 extends beyond missing outcomes to the case of incompletely observed predictors.
Design. We estimate $\beta=(\beta_{0},\beta_{1},\beta_{2})$ in the linear model $y_{i}=\beta_{0}+\beta_{1}x_{1i}+\beta_{2}x_{2i}+\epsilon_{i}$, where $x_{1i}\sim N(0,1)$, $x_{2i}\sim\mathrm{Bernoulli}(0.5)$, $\epsilon_{i}\sim N(0,1)$, and $x_{2i}$ is subject to missingness. Sample size is $N=500$ with M=1, [^0] M=1{,}000 replications. Two PS models govern the missingness of $x_{2i}$: PS1 (MAR), $\mathrm{logit}(\pi_{i})=-1+0.5x_{1i}+0.5y_{i}$; PS2 (MCAR), $\mathrm{logit}(\pi_{i})=-1$. Two OR models generate $y_{i}$:
@@
Results. Table 3 reports bias, standard error, and RMSE ($\times 10$) across all scenarios. The full-sample estimator serves as a benchmark with negligible bias and smallest RMSE. The CC estimator shows substantial bias under MAR (OR1PS1 and OR2PS1), confirming that discarding incomplete cases introduces systematic error when the missingness depends on the observed data.
@@
Across all three settings, the simulations tell a consistent story. When both working models are correct, the proposed estimators recover the efficient doubly robust benchmark. When only one working model is correct, they retain the expected robustness. Most importantly, when the PS model is correct and the OR model is misspecified, ET and HD often improve upon classical AIPW, which is the finite-sample pattern predicted by Corollary 3.